# Robust CATE Estimation Using Novel Ensemble Methods

### v1

@@ Abstract
The estimation of Conditional Average Treatment Effects (CATE) is crucial for understanding the heterogeneity of treatment effects in clinical trials. We evaluate the performance of common methods, including causal forests and various meta-learners, across a diverse set of scenarios revealing that each of the methods fails in one or more of the tested scenarios. Given the inherent uncertainty of the data-generating process in real-life scenarios, the robustness of a CATE estimator to various scenarios is critical for its reliability. To address this limitation of existing methods, we propose two new ensemble methods that integrate multiple estimators to enhance prediction stability and performance - Stacked X-Learner which uses the X-Learner [10] with model stacking for estimating the nuisance functions, and Consensus Based Averaging (CBA), which averages only the models with highest internal agreement. We show that these models achieve good performance across a wide range of scenarios varying in complexity, sample size and structure of the underlying-mechanism, including a biologically driven model for PD-L1 inhibition pathway for cancer treatment.
@@ 1 Introduction
Clinical trials aim to estimate the safety and efficacy of a treatment, usually in comparison to some control. Efficacy is measured by one or more clinical outcomes which are commonly referred to as the trial’s Endpoints (EP). The main measure of interest in terms of efficacy is the Treatment Effect, which is the expected difference in outcome under treatment vs outcome under control. Since for each patient only one potential outcome is observed, a natural way to infer the treatment effect is to use multiple patients - some given the treatment and some the control and compare their outcomes on average. If the assignment of treatment is independent of the potential outcomes (i.e. the probability of being treated is not affected by the expected efficacy of the treatment), then the difference between average outcome of the treated patients and the average outcome of the control patients equals the Average Treatment Effect (ATE) [2].
@@
Where $A$ is the treatment assignment (0 for control, 1 for treatment), and $Y^{a}$ is the outcome under treatment assignment $a$. In clinical trials, the independence of treatment from potential outcomes is achieved by randomization.
@@
Often, there is considerable variability in the treatment effect between different patients [3, 4], making the ATE an insufficient indicator for the treatment’s performance. In such cases, the Conditional Average Treatment Effect (CATE) - the expected treatment effect, conditional on the patient’s features, $X$ is of great interest.
@@
CATE estimation has been intensively studied in recent years in other fields [5, 6, 10, 15], as well as in the clinical field [7]. Two main approaches that have gained popularity are Meta-Learners and Causal Forests. Meta-Learners decompose the problem of estimating the unobserved causal effect into sub-problems of predicting nuisance functions - the different potential outcomes, and the propensity. Meta-Learners can leverage any Machine-Learning (ML) model to estimate the nuisance functions. Several different frameworks were suggested on how to combine these different estimation to obtain a CATE estimator (including S-Learner, T-Learner and X-Learner which are all described in section 3, among others). Causal Forests generalize Random-Forests for the task of CATE estimation by adapting the evaluation and splitting criteria to consider the expected difference in outcomes between treatments instead of the expected outcome. These methods can differ quite substantially from each other and different models have been shown to outperform/under-perform compared to others in different settings [10, 6].
In the analysis of real datasets, the underlying Data-Generation-Process (DGP) is unknown. Furthermore, since CATE cannot be directly observed, standard cross-validation methods frequently used in prediction settings are not applicable. This emphasizes the importance of CATE estimation methods that perform well across various scenarios. To this end we propose two new ensemble methods for CATE estimation - Stacked X-Learner which is an application of the X-Learner Meta-Learner using stacked ML predictors for the nuisance functions, and the Consensus Based Averaging (CBA) which compares several CATE estimates, selects a subset of models that yield the estimates with highest internal agreement and averages them. Both methods are described in detail in section 3. A wider comparison of additional ensemble methods will be explored in future work.
@@
Since our interest is in the analysis of clinical trials data, we designed our evaluation process to emphasise the unique challenges and characteristics of such datasets. In addition to using various additive models based on combinations of linear and nonlinear transformations of features, we also used a biologically driven DGP, which utilises a graphical functional model built to describe the Programmed-Death-Ligand 1 (PD-L1) inhibitor in cancer treatment. A distinct characteristic of clinical trials, especially in Phase II exploratory trials, are their relatively small sample sizes, which make it challenging to learn complex relations from data. Furthermore, it makes it difficult to utilize techniques involving sample splitting. Therefore, this paper focuses on small datasets, ranging from 100 to 750 patients, which represent the setting of interest, as opposed to existing comparative analyses of CATE that consider a higher number of observations [5, 12].
@@
One of the main tasks in the analysis of clinical data is Subgroup identification [9] - the process of identifying a suitable target population for a treatment. We hypothesize that good CATE estimation can be beneficial for guiding the subgroup identification process. In order to keep the focus of this paper on CATE estimation, we leave an in-depth analysis of model-based subgroup search to future work, and only propose a designated evaluation method for CATE estimators, which highlights their accuracy on potential subgroups.
@@
In the next section we introduce the relevant notation and terminology. Section 3 describes in detail the synthetic scenarios used in the analysis, which CATE estimators were compared and the metrics used for the estimators’ evaluation. Results are then presented in section 4. This work focuses on continuous endpoints. Future work will be dedicated to other common endpoint types.
@@ 2 Notation and Terminology
  • $a_{i}$ treatment assignment/allocation. We limit the discussion in this paper to binary treatments, $a_{i}=0$ indicates assignment of patient $i$ to the control and $a_{i}=1$ indicating assignment to treatment.
@@
We follow the framework used by Lipkovich et al. [9] for decomposing $Y$ into prognostic and predictive components (with some changes in notation for internal consistency):
@@
3 Methods
Our study compares the performance of different CATE estimation models in different scenarios. However, the treatment effect is not directly observable in real patients, since only one potential outcome can be measured per patient. Therefore, simulated data is required in order to test the performance in a valid and reliable manner.
We conduct a thorough simulation study on a diverse set of synthetic data generating models. This section reviews in detail the content of the study - the simulated scenarios, the methods tested, and the metrics used to evaluate their performance.
3.1 Data Generation Process
In order to evaluate the robustness of the models, we employed two different types of DGPs. The first type of DGP we employed utilizes mechanistic disease models [13] that form concise descriptions of some main pathway or mechanism of action for the progression of a selected disease. Such models require in depth biological knowledge and are costly to construct, but can generate more realistically distributed data sets with complex interactions formed by the mechanistic relations defined in the model. We utilized a simplified mechanistic model for the PD-L1 pathway in urothelial cancer [14]. This model describes the connections between a few key factors related to the progression of urothelial tumors and specifically to their treatment using PD-L1 inhibitors. The model is described in detail in appendix C.
All DGPs were used to generate training sets of sizes that are representative of clinical trial datasets, ranging from 100 patients to 750 patients. The performance was then evaluated on held-out test sets of 5000 patients.
3.2 Metrics
The evaluation of the different methods’ accuracy employed two metrics, each aimed at measuring a different performance aspects:
  • sRMSE: scaled Root Mean Squared Error
    $sRMSE:=\frac{\sqrt{\sum_{i}(\hat{\tau}_{i}-\tau_{i})^{2}/n}}{\sqrt{\sum_{i}(\tau_{i}-\bar{\tau})^{2}/n}}=\sqrt{\frac{\sum_{i}(\hat{\tau}_{i}-\tau_{i})^{2}}{\sum_{i}(\tau_{i}-\bar{\tau})^{2}}}$
    This is aimed at measuring the distance between predictions and actual CATE. Similar to RMSE, larger errors contribute substantially more to the criteria compared to smaller errors. We scale the RMSE by the standard deviation of the CATE to obtain a scenario invariant metric that is easily interpretable and gives a sense of the relationship between the typical estimation error and the overall dynamic range of the predicted values.
  • RoD: Rate of Discordance
    $RoD:=\frac{1-K(\hat{\tau},\tau)}{2}$
    Where $K$ marks Kendall’s rank correlation coefficient, usually referred to as $\tau_{Kendall}$ (we replace with $K$ here to avoid confusion with the CATE). We call this the rate of discordance because it is equivalent to the proportion of pairs (out of all possible pairs of observations) that are incorrectly ordered by the estimator.
In addition to calculating these metrics over individuals, we also propose an adaptation that averages errors over potential subgroups (SG). We define a set of subgroups of interest (for example - all half-open boxes defined by two variables and having some minimal size) and then compose all such subgroups and treat each subgroup as an effective “observation”, using the following transformations:
$\tau_{i}\mapsto\tau_{SG}:=\frac{\sum_{i\in{SG}}\tau_{i}}{|SG|}$
and
$\hat{\tau}_{i}\mapsto\hat{\tau}_{SG}:=\frac{\sum_{i\in{SG}}y_{i}\cdot a_{i}}{\sum_{i\in{SG}}a_{i}}-\frac{\sum_{i\in{SG}}y_{i}\cdot(1-a_{i})}{\sum_{i\in{SG}}(1-a_{i})}.$ (2)
We then calculate the same accuracy metrics (sRMSE and RoD) using $\tau_{SG}$ and $\hat{\tau}_{SG}$ as the actual and predicted values, accordingly. For subgroups, a natural CATE is observable, unlike the CATE of single observation, which cannot be measured in real life applications. These metrics are more relevant for evaluating performance on a predetermined set of subgroups. CATE estimation methods which perform well on the subgroup metrics can be recommended when the aim of the CATE estimation is to serve as a basis for identifying subgroups.
3.2.1 Subgroup Definition
3.3 Estimators
@@
This analysis uses two general types of estimators - Causal Forests [15] and Meta-Learners [10]. For each, we have explored a wide range of possible configurations with the aim of reaching robust performance across scenarios.
@@
Causal Forest (CF) is an adaptation of random forests - an average of regression trees. In causal forest, the base trees composing the forest are causal trees, aimed at estimating local differences between average potential outcomes [11]. Adapting ideas of doubly robust estimation for CATE [16], CF also has debiased implementations for mitigating potential bias caused by non-random treatment allocation [17].
@@
Meta-Learners are an estimation framework for treatment effect, that enables using any ML model as a “base learner” for learning various nuisance functions and composing an estimator for CATE using a transformation of the learned functions [10, 19]. There are several common meta-learner structures, out of which we compared the following:
@@
  • S learner: A single model is trained to regress the outcomes on the features and the treatment assignment (the treatment is treated as an additional binary variable attached to $X$).
@@
Each of the aforementioned methods can use any ML model as a ”base estimator” for any of the nuisance functions it estimates. The choice of base estimators should account for the assumed complexity of the underlying data-generating process, the size of available training data etc. Clinical trials often test complex mechanisms, which favor more flexible ML models. However, the typical sample sizes in this setting are usually quite limited, favoring simpler more regularized models. Therefore, a set of base-learners spanning a wide range of complexity levels was compared:

### v3

@@ Abstract
The estimation of Conditional Average Treatment Effects (CATE) is crucial for understanding the heterogeneity of treatment effects in clinical trials. We evaluate the performance of common methods, including causal forests and various meta-learners, across a diverse set of scenarios, revealing that each of the methods struggles in one or more of the tested scenarios. Given the inherent uncertainty of the data-generating process in real-life scenarios, the robustness of a CATE estimator to various scenarios is critical for its reliability. To address this limitation of existing methods, we propose two new ensemble methods that integrate multiple estimators to enhance prediction stability and performance - Stacked X-Learner which uses the X-Learner with model stacking for estimating the nuisance functions, and Consensus Based Averaging (CBA), which averages only the models with highest internal agreement. We show that these models achieve good performance across a wide range of scenarios varying in complexity, sample size and structure of the underlying-mechanism, including a biologically driven model for PD-L1 inhibition pathway for cancer treatment. Furthermore, we demonstrate improved performance by the Stacked X-Learner also when comparing to other ensemble methods, including R-Stacking, Causal-Stacking and others.
@@ 1 Introduction
Clinical trials aim to estimate the safety and efficacy of a treatment, typically in comparison to some control group. Efficacy is measured by one or more clinical outcomes, commonly referred to as the trial’s Endpoints (EP). The primary measure of interest in terms of efficacy is the Treatment Effect, which is the expected difference in outcome between the treated patients and the control patients. Since only one potential outcome is observed for each patient, the treatment effect can be inferred by comparing the average outcomes of multiple patients—some given the treatment and some the control. If treatment assignment is independent of potential outcomes (i.e., the probability of treatment is unaffected by the expected treatment efficacy), the difference between the average outcomes of treated and control patients represents the Average Treatment Effect (ATE) Rubin (1974).
@@
where $A$ is the treatment assignment (0 for control, 1 for treatment) and $Y^{a}$ is the outcome under treatment assignment $a$. In clinical trials, the independence of treatment from potential outcomes is achieved by randomization.
@@
There is often considerable variability in treatment effects among different patients McCormick et al. (2003); Kravitz et al. (2004), making ATE an insufficient indicator of treatment performance. In such cases, the Conditional Average Treatment Effect (CATE) - the expected treatment effect, conditional on the patient’s features, $X$ is of great interest. Other parameters of the individual treatment effect distribution can be of interest as well, e.g., the proportion of treatment responders (Frostig et al., 2024).
@@
In recent years, CATE estimation has been extensively studied in various fields (Curth et al., 2021; Künzel et al., 2019; Wager and Athey, 2018; Jacob, 2021), including the clinical field (Curth et al., 2024). Two main approaches that have gained popularity are Meta-Learners and Causal Forests. Meta-Learners decompose the problem of estimating the unobserved causal effect into sub-problems of predicting nuisance functions: the different potential outcomes and the propensity. Meta-Learners can leverage any Machine-Learning (ML) model (called base-learners) to estimate the nuisance functions. Several different frameworks were suggested on how to combine these different estimates to obtain a CATE estimator (including S-Learner, T-Learner, X-Learner and R-Learner which are all described in section 3, among others). Causal Forests generalize Random-Forests for the task of CATE estimation by adapting the evaluation and splitting criteria to consider the expected difference in outcomes between treatments instead of the expected outcome.
In real-life applications, the underlying Data Generation Process (DGP) is unknown. Additionally, because CATE cannot be directly observed, standard cross-validation methods used in supervised learning settings are inapplicable. This challenge has led to numerous studies on how to select the best CATE estimator by examining various loss functions.
Schuler et al. (2018) found that a loss function based on the R-Learner performs best for CATE selection. Mahajan et al. (2022) extended this work, finding that with an Auto-ML approach to the base-learners, the T-Learner and X-Learner based losses outperform the R-Learner based loss. Doutreligne and Varoquaux (2023) recommended using the R-Loss for model selection, stacking for nuisance function estimation, and a 90/10% split for model estimation and selection. Curth et al. (2024) compared selection metrics based on plug-in estimates, factual estimation, and pseudo-outcomes, arguing that no single metric is universally recommended as it depends on the dataset. However, they recommend the pseudo-outcome as the default approach.
An alternative approach to selecting the best CATE estimator involves averaging multiple CATE estimators rather than choosing a single one. Nie and Wager (2021) proposed R-Stacking, which combines several CATE estimators by finding the optimal linear combination that minimizes the R-Loss, leveraging the strengths of each individual estimator. Han and Wu (2022) introduced Causal-Stacking for randomized experiments, obtaining a linear combination of CATE estimators that minimizes the mean squared error of a pseudo outcome based on the , with non-negative weights that sum to one. It was shown to outperform R-Stacking in balanced treatment allocation settings.
@@
In this work our goal is to study CATE estimation using ensembles in clinical trial data. We designed our evaluation process to emphasise the unique challenges and characteristics of such datasets. In addition to using various additive models based on combinations of linear and nonlinear transformations of features, we also used a biologically driven DGP, which utilises a graphical functional model built to describe the Programmed-Death-Ligand 1 (PD-L1) inhibitor in cancer treatment. A distinct characteristic of clinical trials, especially in Phase II exploratory trials, is their relatively small sample sizes, which makes it challenging to learn complex relations from data. Furthermore, it makes it difficult to utilize techniques involving sample splitting. Therefore, we focus on small datasets, ranging from 100 to 750 patients, representing the setting of interest. Current benchmarks of CATE estimation usually consider a higher number of observations (Schuler et al., 2018; Jacob, 2021; Mahajan et al., 2022; Acharki et al., 2023; Doutreligne and Varoquaux, 2023). Only Nie and Wager (2021); Curth et al. (2024) considered data-sets with 500 observations.
Furthermore, we propose two new ensemble methods for CATE estimation - Stacked X-Learner which is an application of the X-Learner Meta-Learner using stacked ML predictors for the nuisance functions, and the Consensus Based Averaging (CBA) which compares several CATE estimates, selects a subset of models that yield the estimates with highest internal agreement and averages them. Both methods are described in section 4.
@@
One of the main tasks in the analysis of clinical data is Subgroup identification Lipkovich et al. (2017) - the process of identifying a suitable target population for a treatment. We hypothesize that good CATE estimation can be beneficial for guiding the subgroup identification process. In order to keep the focus of this paper on CATE estimation, we leave an in-depth analysis of model-based subgroup search to future work, and only propose a designated evaluation method for CATE estimators, which highlights their accuracy on potential subgroups.
@@
In the next section we introduce the relevant notation and terminology. Section 3 overviews different CATE estimators and ensembles of CATE estimators. In section 4 we propose our two ensemble methods. In section 5 we describe the data-generation-process (DGP) used in the simulation, section 6 is devoted to the evaluation metrics. In section 7 we present our simulation study and results. Future work will be dedicated to other common endpoint types.
@@ 2 Notation and Terminology
  • $A_{i},a_{i}$ treatment assignment/allocation. We limit the discussion in this paper to binary treatments, $a_{i}=0$ indicates assignment of patient $i$ to the control and $a_{i}=1$ indicating assignment to treatment.
  • $\pi(x_{i})$ the propensity score of each patient assuming strong ignorability, then $\pi(x_{i})=\mathbb{P}(a_{i}=1)$.
@@
We denote estimators using $\hat{\quad}$, e.g., a CATE estimator is denoted by $\hat{\tau}$. We follow the framework used by Lipkovich et al. (2017) for decomposing $Y$ into prognostic and predictive components (with some changes in notation for internal consistency):
@@
3 Related Work
@@
This analysis uses two general types of estimators - Causal Forests (Athey and Imbens, 2016; Wager and Athey, 2018) and Meta-Learners (Künzel et al., 2019). For each, we have explored a wide range of possible configurations with the aim of reaching robust performance across scenarios.
@@
Causal Forest (CF) is an adaptation of random forests - an average of regression trees. In causal forest, the base trees composing the forest are causal trees, aimed at estimating local differences between average potential outcomes Athey and Imbens (2016). To handle issues of selective inference they introduced the notion of honesty, where the sample is split into a training and estimation set. The training set is used to identify the partitions of the tree and the estimation set to estimate the CATE within the partition. Furthermore, adapting ideas of doubly robust estimation for CATE Chernozhukov et al. (2017), CF also has debiased implementations for mitigating potential bias caused by non-random treatment allocation Wager and Athey (2018).
@@
Meta-Learners are an estimation framework for treatment effect, that enables using any ML model as a “base learner” for learning various nuisance functions and composing an estimator for CATE using a transformation of the learned functions Künzel et al. (2019); Kennedy (2023). There are several common meta-learner structures, out of which we compared the following:
@@
  • S learner: A single model is trained to regress the outcomes on the features and the treatment assignment (the treatment is modeled as an additional binary variable attached to $X$).
@@
  • R learner: This approach suggests to minimize the following loss function,
    $\tau(\cdot)=\arg\min_{\tau}\left\{E\left(\left[\{Y_{i}-m(X)\}-\{A_{i}-\pi(X)\}\tau(X)\right]^{2}\right)\right\},$
    where $m(X_{i})=\mathbb{E}(Y_{i}|X_{i})=\mu_{0}(X_{i})+\pi(X)\tau(X)$. In practice, the nuisance functions $m(X),\pi(X)$ are estimated in a cross-fitting manner.
@@
Each of the aforementioned methods can use any ML model as a ”base learner” for any of the nuisance functions it estimates. The choice of base estimators should account for the assumed complexity of the underlying data-generating process, the size of available training data etc. Clinical trials often test complex mechanisms, which favor more flexible ML models. However, the typical sample sizes in this setting are usually quite limited, favoring simpler more regularized models.
3.1 Ensemble Methods
We will assume $K$ CATE estimation models denoted by $\hat{\tau}_{1},\ldots,\hat{\tau}_{K}$. Our goal is to find a function of the estimators which results in a single ensemble estimator, denoted by $\hat{\tau}^{*}$.
  • R-Stacking: R-Stacking is based on finding the best linear combination of $\hat{\tau}_{1},\ldots,\hat{\tau}_{K}$ which minimizes the R-loss (Nie and Wager, 2021),
    $(\hat{b},\hat{c},\hat{\alpha})=\arg\min_{b,c,\alpha}\left\{\sum_{i=1}^{n}\left[\left\{Y_{i}-\hat{m}(X_{i})\right\}-b-\left\{c+\sum_{k=1}^{K}\alpha_{k}\hat{\tau}_{k}(X_{i})\right\}\left\{A_{i}-\hat{\pi}(X_{i})\right\}\right]^{2}\right\},$
    subject to $\alpha_{k}>0,\forall k$. $n$ is the number of observation in the validation set. The CATE estimation according to the ensemble is
    $\hat{\tau}^{*}(x)=\hat{c}+\sum_{k=1}^{K}\alpha_{k}\hat{\tau}_{k}(x).$
  • Causal Stacking: Han and Wu (2022), focused on the setting of a randomized experiment, i.e., the true propensity, $\pi(x)$ is known. Their suggestion is to obtain a linear combination of estimators which minimizes the mean square error of a pseudo outcome,
    $\hat{\alpha}=\arg\min_{\alpha}\left\{\sum_{i=1}^{n}\left[\hat{\tau}_{i}^{\mathrm{PO}}-\sum_{k=1}^{K}\alpha_{k}\hat{\tau}_{k}(X_{i})\right]^{2}\right\},$
    subject to $\sum_{k=1}^{K}\alpha_{k}=1$ and $\alpha_{k}>0,\forall k$. The psuedo outcome is defined by,
    $\hat{\tau}_{i}^{\mathrm{PO}}=\left(\hat{\mu}_{1}(X_{i})-\hat{\mu}_{0}(X_{i})\right)+\frac{\left(Y_{i}-\hat{\mu}_{1}(X_{i})\right)A_{i}}{{\pi}(X_{i})}-\frac{\left(Y_{i}-\hat{\mu}_{0}(X_{i})\right)\left(1-A_{i}\right)}{1-{\pi}(X_{i})}.$
    Causal-Stacking was shown to perform better than the R-Stacking when the treatment allocation is balanced. Both methods require that the estimated nuisance functions, $\hat{m},\hat{\pi},\hat{\tau_{k}}$ or $\hat{\mu}_{1},\hat{\mu}_{0}$ are estimated on a different set then the one used for obtaining $\hat{\alpha}_{k}$
  • T-Stacking: The ideal method to train an ensemble, would be to regress the resulting CATE models estimates against the true CATE of a validation set. Unfortunately, we can only estimate CATE and not directly observe it. To circumvent the issue, it was suggested to replace the true CATE with some CATE estimator trained on the validation set. The T-score, is one such loss, where the CATE is estimated using a T-Learner (Alaa and Van Der Schaar, 2019; Mahajan et al., 2022). T-Stacking is an adaptation of the T-score for the purposes of ensembeling,
    $\hat{\alpha}=\arg\min_{c,\alpha}\left\{\sum_{i=1}^{n}\left[\hat{\tau}_{T}(X_{i})-\sum_{k=1}^{K}\alpha_{k}\hat{\tau}_{k}\right]^{2}\right\}$
    subject to $\alpha_{k}>0$. Here $\hat{\tau}_{T}$ is a T-Learner trained on the validation-set.
4 Suggested Ensembles
We suggest two novel methods of ensembling. One, motivated by the work of Doutreligne and Varoquaux (2023), is based on the X-Learner and stacking for the estimation of the nuisance functions. The second method is based on finding consensus among models and does not require sample splitting to obtain the ensemble estimator. This method is motivated by the need to efficiently use the limited sample size available in clinical trials.
  • Stacked X-Learner: Motivated by the results of Doutreligne and Varoquaux (2023), we suggest an X-Learner based on stacking. We use stacking for the first level ($\mu_{0}$, $\mu_{1}$) and second level ($\tau_{0}$, $\tau_{1}$) of nuisance functions. Each function is estimated as a standard supervised learning task, making it feasible to use the typical stacking methods.
  • Consensus Based Averaging (CBA): Existing ensemble methods are not sample efficient; they require splitting the data into a training set, where the nuisance functions and the CATE are estimated, and a validation (or averaging) set, which is used to train the ensemble estimator. This method prioritizes models with higher agreement. Initially, predictions from each model are collected. Kendall’s Tau correlation coefficients are calculated between each pair of models, and the mean correlation coefficient for each model is determined. Models are then sorted based on their mean correlation values. The method identifies a subset of models that are in higher agreement by detecting the ’knee’ in the sorted mean correlation curve. Models up to this knee point are selected, and their predictions are averaged to produce the final ensemble prediction, thereby enhancing robustness and accuracy. For further details on the method, see appendix A.
5 Data Generation Process
We considered three groups of DGP. The first uses a functional graphical model (Scholz et al., 2013) to concisely model the PD-L1 pathway in urothelial cancer. The second is a flexible additive linear model with nonlinear transformations and second order interactions, which allows to easily create diverse and tunable settings. The third DGP is based on the ACIC2016 datasets Dorie et al. (2019), which use a combination of real-world variable data and synthetic specified treatment effects.
5.1 PD-L1-biological data generation model
This DGP aims to mimic a biological process related to the progression of tumors, a setting often encountered in clinical trials. The DGP employs a mechanistic disease model that forms a concise description of a main pathway or mechanism of action for the progression of a selected disease. Such models require in-depth biological knowledge and are costly to construct, but they can generate more realistically distributed datasets with complex interactions formed by the mechanistic relations defined in the model. We utilized a simplified mechanistic model for the PD-L1 pathway in urothelial cancer (Mariathasan et al., 2018). This model describes the connections between a few key factors related to the progression of urothelial tumors and specifically to their treatment using PD-L1 inhibitors.
  • Tumor Growth: Tumor growth is driven by the mutation burden (increasing the growth), while it is inhibited by the post-treatment levels of CD8+ effector cells.
[…]
5.2 Linear Models
5.3 ACIC 2016 Datasets
The ACIC 2016 datasets consist of real-world covariates, with simulated treatment and response. Using a mixture of real-world and simulated data provides an appealing compromise between maintaining the realistic complex structure of clinical data, and still having full knowledge of potential outcomes and their expected values.
The covariates are taken from a longitudinal study aimed at identifying the causal covariates responsible for developmental issues Dorie et al. (2019). From the large dataset, 4802 observations and 58 covariates were selected. Of these covariates, 3 are categorical, 5 are binary, 27 are count data, and the remaining 23 are continuous.
Several datasets are created using the covariates, representing a large variety of treatment allocation mechanisms, prognostic and predictive factors Dorie et al. (2019). These datasets are frequently referenced and used in other comparative analyses, e.g. Mahajan et al. (2022); Han and Wu (2022); Curth and Van Der Schaar (2023). Focusing on clinical trial data, we sub-sampled the dataset into smaller cohorts and used the remainder as the models’ evaluation dataset. Furthermore, we have changed the treatment allocation to be balanced and randomized.
6 Metrics
The evaluation of the different methods’ accuracy employed two metrics, aimed at measuring different performance aspects:
  • sRMSE: scaled Root Mean Squared Error. The sRMSE is calculated as follows:
    $\mathrm{sRMSE}:=\frac{\sqrt{\sum_{i}(\hat{\tau}_{i}-\tau_{i})^{2}/n}}{\sqrt{\sum_{i}(\tau_{i}-\bar{\tau})^{2}/n}}=\sqrt{\frac{\sum_{i}(\hat{\tau}_{i}-\tau_{i})^{2}}{\sum_{i}(\tau_{i}-\bar{\tau})^{2}}}$
    sRMSE measures the distance between predictions and actual CATE. Similar to RMSE, larger errors contribute substantially more to the criteria compared to smaller errors. We scale the RMSE by the standard deviation of the CATE to obtain a scenario invariant metric that is easily interpretable and gives a sense of the relationship between the typical estimation error and the overall dynamic range of the predicted values.
  • RoD: Rate of Discordance
    $\mathrm{RoD}:=\frac{1-\mathcal{K}(\hat{\tau},\tau)}{2}$
    Where $\mathcal{K}$ marks Kendall’s rank correlation coefficient.  [^1] Kendall Tau is usually referred to as $\tau$. We denote it by $\mathcal{K}$ here to avoid confusion with the CATE. We call this the rate of discordance because it is equivalent to the proportion of pairs (out of all possible pairs of observations) that are incorrectly ordered by the estimator.
In addition to calculating these metrics over individuals, we also propose an adaptation that averages errors over potential subgroups (SG). We denote a subgroup by $g_{j}$ which is a set of patients indices. We define $\{g_{1},\ldots,g_{G}\}$, the set of all potential subgroups of interest (for example - all half-open boxes defined by two variables and having some minimal size). For each subgroup we compute the true and estimated CATE,
$\tau_{g_{j}}:=\frac{\sum_{i\in{g_{j}}}\tau_{i}}{|g_{j}|}$
and
$\hat{\tau}_{g_{j}}:=\frac{\sum_{i\in{g_{j}}}y_{i}\cdot a_{i}}{\sum_{i\in{g_{j}}}a_{i}}-\frac{\sum_{i\in{g_{j}}}y_{i}\cdot(1-a_{i})}{\sum_{i\in{g_{j}}}(1-a_{i})}.$ (2)
We then calculate the same accuracy metrics (sRMSE and RoD) using $\tau_{SG}$ and $\hat{\tau}_{SG}$ as the actual and predicted values, accordingly. E.g.,
$\mathrm{sRMSE}_{\mathrm{SG}}=\sqrt{\frac{\sum_{j=1}^{G}(\hat{\tau}_{g_{j}}-\tau_{g_{j}})^{2}}{\sum_{j=1}^{G}(\tau_{g_{j}}-\bar{\tau}_{\mathrm{SG}})^{2}}},$
where $\bar{\tau}_{\mathrm{SG}}=\frac{1}{G}\sum_{j=1}^{G}\tau_{g_{j}}$. These metrics are more relevant for evaluating performance on a predetermined set of subgroups. CATE estimation methods which perform well on the subgroup metrics can be recommended when the aim of the CATE estimation is to serve as a basis for identifying subgroups.
6.1 Subgroup Definition
7 Simulation Study
Our study compares the performance of various CATE estimation models in different scenarios. The treatment effect is not directly observable in real patients, since only one potential outcome can be observed per patient. Therefore, simulated data is required in order to test the performance in a valid and reliable manner.
7.1 Methods
We consider a variety of Meta-Learners, including the S-Learner, X-Learner, T-Learner, and DR-Learner. The performance of Meta-Learners is strongly influenced by the base-learners used. We included a set of base-learners spanning a wide range of complexity levels: